• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 23:45:27 on October 5, 1998

    Entry number 2482

    keyword=Beam Time Accounting
    Hall A

    10/5/1998 1600-2400
    Experiment #: E94-010
    Shift Leader: prout
    Shift Workers: McCormick, Zolnierczuk, Jiang

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    1 1 1 1 30 30 1 0 6
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    0 0 0 0 30 30 0 0 1
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 1 1
    Experiment
    Ready
    1 1 1 1 1 1 1 1 8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 16:00 Moller measurement begins. An NMR measurement was made also. 18:20 Moller complete (pol = 67%) Start NMR while Moller target is moved out. 18:40 Problem with Coda : no data file was created when run1610 was started. Bob says this is the same problem they had over the weekend. We restart Coda and it works again. 19:15 Finally running after readjustment of beam spot (7mm too high) 20:30 Down because of RF cavity problem 21:30 Problem fixed start data taking again 22:30 Stop to do NMR and to do current calibration Have confirmed half wave plate was inserted 23:00 MCC takes over beam for maintanence

    Data Entry by: Prout